4. If tan A = cot B, prove that A + B = 90°.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
tan A = cot B We know that cot B = tan (90° – B) To prove A + B = 90°, substitute the above equation in the given problem tan A = tan (90° – B) A = 90° – B A + B = 90° Hence Proved.